OpenStudy (anonymous):

Solve the equation for X x + 7 = 13

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

That's what I was thinking (:

OpenStudy (anonymous):

x+7=13 x+7-7=13-7 x=6

OpenStudy (anonymous):

I got it by going x+7=13 13-7=6 so 6+7=13 so x=6